# Remove point-mass concept - remove singularities from GR

### v2

@@ III.5 Determination of $\kappa$ Tully- Fisher relation
  • Beyond the visible disks of galaxies, the observed $v^{2}$ has , more often than not, the flat asymptote $v^{2}\propto\sqrt{M}$, rather than $\propto M$. See Sobouti:2008xz ; shostak1973aperture ; Bosma:1981zz ; Begeman:1989kf ; Begeman:1991iy ; Sanders:1998gr ; Sanders:2002pf ; Persic:1995tc ; Persic:1995ru .
@@
Milgrom is silent as to whether his $a_{0}$ is a universal acceleration applicable to all spirals or not. There are, however, indications that $a_{0}$ might depend on the size and/or mass of the host galaxies, see Sobouti:2006rd for $a_{0}$ in a list of 53 spiral galaxies.
@@ III.6 Rotation curves
$v^{2}=\frac{1}{2}c^{2}r\frac{dB}{dr},$
where, for the far zone $\frac{1}{2}c^{2}\frac{dB}{dr}$ is spelled out in (31). Fig.(1) is the rotation curve for our toy galaxy. The curve consists of three patches pieced together. Its free parameter are $N=10^{70},\,\,\,\kappa=10^{-3},\,\,\,k=0.2,\,\,\,f=0.001,\,\,\,\alpha=0$. The model has a central blackholes of $0.001$th of the galaxy. Superimposed on the almost flat branch of the curve, there are short-wavelength and small-amplitude sinusoidal variations. They come from the degenerate and phase-tuned matter of the central blackhole. Compare them with similar variations in Fig.(3) for the rotation curves of M31 and NGC 6503. To the best of our knowledge they are never before, addressed in the astronomical literature. See also Fig.(2), for assorted rotation curves from Carroll and Ostlie, Carroll-17 ; Wilson et al., Rholf ; Freedman and KaufmannKaufmann .
IV Massive Gravitons
A Historical Note: Yukawa introduced his field in 1935 to explain the residual strong interactions that bind nucleons together within a nucleus at femtometer scales. The mediating $\pi^{\pm}$ were discovered in 1947 and $\pi^{0}$ in 1950. There are two masses associated with Yukawa paradigm: a) the masses of interacting quarks between nucleons and b) the masses of the mediating pions.
Our YKG serves is the gravitational analog of the Yukawa field, albeit operating on galactic scales. There are also two masses in YKG: a) the mass of the gravitating nucleons incorporated in $k=r_{s}\frac{mc}{h}=\frac{2Gm^{2}}{ch}$, (12); and b) the mass of the mediating graviton. The Compton mass and wavelength associated with $\kappa$ are,
@@ IV Massive Gravitons
$m_{g}=\frac{\hbar}{c\lambda_{g}}=1.16\times 10^{-29}\left(\frac{M}{m}\right)^{-1/2}~{}\textrm{kg}=6.49\times 10^{4}\left(\frac{M}{m}\right)^{-1/2}\textrm{eV}\,.$ (34)
@@
Two Features to Note: a) Transition from the Newtonian to the non-Newtonian regime takes place, where the two forces become equal,
@@
Mass (kg) Mass ($\sqrt{\textrm{Mass (kg)}}$ $\lambda_{g}(m)$ $m_{g}$
MW $1.5\times 10^{42}$ $1.22\times 10^{21}$ $1.2\times 10^{21}\approx 38~{}\textrm{kpc}$ $9.5\times 10^{-29}$eV
Sun $2\times 10^{30}$ $1.4\times 10^{15}$ $1.06\times 10^{15}\approx 7100~{}\textrm{AU}$ $1.7\times 10^{-22}$ eV
Earth $6\times 10^{24}$ $2.44\times 10^{12}$ $1.7\times 10^{12}\approx 11.3~{}\textrm{AU}$ $1.06\times 10^{-19}$eV
1 kg $1$ 1 $0.74$ $26\times 10^{-6}$ev
@@
[…]
Table 1: Compton wavelength and mass of massive gravitons that mediate the YKG gravitation of a wide range of masses, from celestial bodies to quarks.
In Table (1) we have calculated the mass and the wavelength of gravitons for the Milky Way, the Sun, the Earth, one kg mass, nucleons, and $u$ and $d$ quarks
@@
  • For the Sun (or rather the whole solar system) transition takes place at $7100$ AU, somewhere within the Oort clouds, believed to have an inner edge of $2000$ to $5000$ AU and an outer edge of $10000$ to $100000$ AU, see https://science.nasa.gov/solar-system/oort-cloud/.
@@
  • The last four entries for $p,n,u,d$ are included as a matter of curiosity; and they should not be taken seriously at this stage. One, however, wonders how and why the gravitons mediating the gravitational interaction of $u$ and $d$ and pions mediating the residual strong interaction of a smae $u$ and $d$ have almost the identical masses, $\simeq 130-140~{}\textrm{MeV}$. There must be a deeper connection between the YKG gravity and, at least Yukawa’s residual, strong interactions.
There is a rich literature on the mass of graviton depending on how the assumed gravitation differs from the standard GR (Isham:1971gm, ; Deser:1981wh, ; deRham:2010kj, ; deRham:2014zqa, ; Arkani-Hamed:2002bjr, ), or how they are deduced from astronomical observations. Closest to what we propose here is Desai:2018swo . There a Yukawa-ameliorated gravitation is used to deduce the mass of graviton from observation on Abell1689 galaxy cluster. They report
@@
This graviton mass is a factor of about $7$ lighter than our $9.5\times 10^{-29}$ eV for the Milky Way. This means that the graviton from Abell 1689 reaching earth-bound observers, has trespassed a mass of $497^{2}=$ 49 Milky Ways. Abell 1689 galaxy cluster is reported to have about 1000 galaxies, Abell:1989mu .
A recent paper, DAgostino:2024ojs , employs a Yukawa cosmology to constrain the graviton mass from observations of the Milky Way and M31. Although their approach and reasoning differs from ours, the fact both they and we speak of Yukawa in galactic and cosmic contexts, conclusions are very much the same. For instance, both studies address deduction of graviton mass from observations of rotation curves, or both maintain rotation curves should be considered as quantum manifestations of galactic matters. Their graviton mass $m_{g}\approx 1.54\times 10^{-62}$ kg $=8.64\times 10^{-27}$ eV is, however, is $91$ times larger than our, $9.5\times 10^{-29}$ eV for the Milky Way.
V Conclusion
We argue that matter in classical physics, including GR, is a collection of discrete point-like entities, to which in defiance of Heisenberg’s principle of uncertainty, one assigns precise coordinates and momenta. Singularities and divergences encountered in Schwarzschild’s solution or elsewhere in GR stem from this non-quantum act. We propose to replace that discrete collection of point-like masses by a collection of wave-like entities, KG and/or YKG waves. On the one hand, waves have spatial distributions and do not end up in spacetime singularities. On the other hand, YKG waves reach infinity and produce $1/r$ forces. This feature alone explains the flat rotation curves of spiral galaxies (two birds with one stone!).
The far zone solutions are the highlight of this paper. YKG waves reach infinity as $\frac{1}{r}\exp(-\kappa/r)\cos(kr+\alpha)$ if degenerate, and as $\frac{1}{r}\exp(-\kappa/r)$ if non-degenerate, see (23), (25), (26) . At intermediate distances, the gravitational force is dominantly $\frac{1}{r^{2}}$, but transits to $\frac{1}{r}$ as $r\rightarrow\infty$. The latter is responsible for the flat rotation curves of spirals and can be interpreted as an alternative gravity, a dark matter scenario, or whatever.
@@ V Conclusion
Nowadays quantum technologists are sending and receiving quantum messages across distances of tens and hundreds of kilometers through the science and technology of quantum entanglement. Do astronomers have sufficient evidence to claim they are observing quantum effects across galaxies through rotation curves? We think they do; the Tully-Fisher relation is the cumulative quantum relic of some $10^{70}$ nucleons in galactic matters.
Thinking ahead: In the standard GR gravitation is mediated by massless gravitons, known to come in two polarizations. Our YKG gravitons are massive and are expected to have at least three polarizations. If so, they are then expected to be in one-to-one correspondence with the three Yukawa pions that mediate the residual strong interactions. There are further similarities between YKG and the residual strong interactions; for example, both are attractive forces; or as shown in Table (1), the mass of YKG graviton, for $u$ and $d$ quarks are nearly identical to the masses of the $\pi$ mesons. Given these numerous commonalities, one might wonder whether YKG gravitation and residual strong interactions are, in fact, two manifestations of a single underlying concept—two sides of the same coin.
Last but not least. In his 2023 paper (Kerr:2023rpn, ), entitled ’Do Black Holes have Singularities?’, Roy Kerr questions Penrose’s singularity theorem and concludes that:
The author’s opinion is that gravitational clumping leads inevitably to black holes, confirming what is observed, but this does not lead to singularities. … These all assume that the matter is classical and that it satisfies whatever nineteenth century equation of state proponents require … The author has no doubt, and never did, that when Relativity and Quantum Mechanics are melded it will be shown that there are no singularities anywhere.
Evidently, Relativity and Quantum Mechanics are not melded yet. There are, however, promising signs at the horizon. If acceptable to Roy, the present paper and those of Desai:2018swo and (DAgostino:2024ojs, ) could be adequate cases. To these we may add the ideas of Novikov (Novikov:2016fzd, ) who advocates a quantum cosmology of an early and inflation-free Universe.
Acknowledgements.
We thank M. Vahidinia and V. Taghiloo for their participation in formulating the physics of the model elaborated in Sec.(II). We also thank H. Firouzjahi for his insightful comments on the Astrophysical aspects of the paper.

### v3

@@ 3.5 Determination of $\kappa$ Tully- Fisher relation
  • Beyond the visible disks of galaxies, the observed $v^{2}$ has , more often than not, the flat asymptote $v^{2}\propto\sqrt{M}$, rather than $\propto M$. See [19, 20, 21, 22, 23, 24, 25, 26, 27].
@@
Milgrom is silent as to whether his $a_{0}$ is a universal acceleration applicable to all spirals or not. There are, however, indications that $a_{0}$ might depend on the size and/or mass of the host galaxies, see [28] for $a_{0}$ in a list of 53 spiral galaxies.
@@ 3.6 Rotation curves
$v^{2}=rg(r)={e^{-2\kappa r/r_{s}}}\left(\frac{{GNm}}{{{r}}}+{{N{c^{2}}\kappa}}\Big{(}1+\frac{3}{2}f\cos 2(kr/r_{s}+\alpha)\Big{)}\right).$
Fig.(1) is the rotation curve of our toy galaxy. We consider a spherically symmetric collection of YKG gravitating waves in its monopole approximation. There are no galactic planes or spiral arms in the model to mimic any realistic spiral or non-spiral galaxy. All we want to demonstrate is that moving outward from the center, there is steep rise to a maximum followed by a gentle declines towards an almost flat asymptote.
The free parameters of the model are $N=10^{70},\,\kappa=10^{-3},\,k=0.2,\,f=0.001\,,\alpha=0.$ The model has a central blackholes of $0.001$th of the galaxy. Superimposed on the almost flat branch of the curve, there are short-wavelength and small-amplitude sinusoidal variations. They come from the degenerate and phase-tuned matter of the central blackhole. Compare them with similar variations in Fig.(2) for the rotation curves of M31 and NGC 6503. To the best of our knowledge they are never before, addressed in the astronomical literature. See also Fig.(3), for assorted rotation curves from Carroll and Ostlie [29]; Wilson et al. [30]; Freedman and Kaufmann [31].
4 Massive Gravitons
A Historical Note: Yukawa introduced his field in 1935 to explain the residual strong interactions that bind nucleons together within a nucleus at femtometer scales. The mediating bosons, $\pi^{\pm}$, were discovered in 1947 and $\pi^{0}$ in 1950. There are two masses associated with Yukawa paradigm: a) the masses of interacting quarks between nucleons and b) the masses of the mediating pions.
Our YKG is the gravitational analog of the Yukawa field, albeit operating on galactic scales. There are also two masses in YKG: a) the mass of the gravitating nucleons incorporated in $k=r_{s}\frac{mc}{\hbar}=\frac{2Gm^{2}}{c\hbar}$, (14); and b) the mass of the mediating graviton. The Compton mass and wavelength associated with $\kappa$ are,
@@ 4 Massive Gravitons
$m_{g}(\lambda_{g})=\frac{\hbar}{c\lambda_{g}}=1.16\times 10^{-29}\left(\frac{M}{m}\right)^{-1/2}~{}\textrm{kg}=6.49\times 10^{4}\left(\frac{M}{m}\right)^{-1/2}\textrm{eV}\,.$ (40)
@@
Two Features to note:
a) Transition from the Newtonian to the non-Newtonian regime takes place, where the two forces become equal,
@@
The $m_{g}(r_{tr})$ of (40) mediates the YKG gravitation at the location $r_{tr}=\lambda_{g}.$ One expects as one moves from $r_{tr}$ to a different location $r$ this mass to change to $m_{g}(r),$ but how? To answer the question we note that the $m_{g}(r_{tr})$ does not have kinetic energy, for the YKG gravitation is static. It does not have a rest mass energy either. There only remains to be a potential energy at $r_{tr}$. To see this, let us multiply $m_{g}(r_{tr})$ by $e^{-2\kappa r_{tr}/r_{s}}$ and divide by the same
$m_{g}(r_{tr})=\frac{\hbar}{cr_{tr}}e^{-2\kappa r_{tr}/r_{s}}e^{2\kappa r_{tr}/r_{s}}.$
We recognize that ${\hbar}/{cr_{tr}}e^{-2\kappa r_{tr}/r_{s}}$ is proportional to the gravitational potential at $r_{tr}$ and conclude that upon moving from $r_{tr}$ to a generic $r$ to have
$m_{g}(r)=\frac{\hbar}{cr}e^{-2\kappa(r-r_{tr})/r_{s}}.$ (44)
We will come back to this (44), when we discuss the graviton deduced from the orbital precession of $S2$ star orbiting the supermassive blackhole in the Constellation Sagittarius A* (SMBH SgrA*) .
In Table (4) we have calculated the mass and the wavelength of gravitons for the Milky Way, the supper massive blackhole Sgr A*, the Sun, the Earth, one kg mass, nucleons, and $d$ and $u$ quarks
@@
  • The case of SMBH Sgr A* will be discussed below.
  • For the Sun (or rather the whole solar system) transition takes place at $7100$ AU, somewhere within the Oort clouds, believed to have an inner edge of $2000$ to $5000$ AU and an outer edge of $10000$ to $100000$ AU, see https://science.nasa.gov/solar-system/oort-cloud.
@@
  • The last four entries for $p,n,d,u$ are included as a matter of curiosity. They should not be taken seriously at this stage. One, however, wonders how and why the gravitons mediating the gravitational interaction of $u$ and $d$, and pions mediating the residual strong interaction of the same $u$ and $d$ have almost identical masses, $\simeq 130-140~{}\textrm{MeV}$. There might be a deeper connection between the YKG gravity and, at least Yukawa’s residual, strong interactions.
The SMBH SgrA* at the center of the Milky Way has a mass of $4.3\times 10^{6}M_{\odot}=8.6\times 10^{36}kg$. The S2 star, a bright infrared source of radiation, orbits SMBH SgrA* with the eccentricity $e=0.885$ and semi major axis $970AU$. There are almost 30 years of observational data on the S2 orbit. In [33, 32] the authors use a Yukawa modified gravity of their own, calculate the periastron precession of the orbit and report a graviton mass of $m_{g}(S2)\leq 2.9\times 10^{-21}eV$.
To compare this graviton mass with the one reported in Table 4 ($8\times 10^{-26}$ eV) let us note the former is obtained from the analysis of the S2 orbit. The rough distance of S2 from SMBH SgrA* is about $\approx 700AU$, the sum of the semi major + semi minor axes divided by 2. The graviton mass we report belongs to a distance $r_{tr}=\lambda_{g}=1.47\times 10^{7}AU$. Considering (44), we have This number is indeed in excellent agreement with $m_{g}(S2)\leq 2.9\times 10^{-21}eV$ reported in [33, 32].
In addition to the massive graviton of SMBH SgrA*, there are further literature on the mass of graviton depending on how the assumed gravitation differs from the standard GR [34, 35, 36, 37, 38], or how they are deduced from astronomical observations. The case of SMBH SgrA* was just discussed and we found in good agreement with what we find. Another closest to what we propose here is [39]. There a Yukawa-ameliorated gravitation is used to deduce the mass of graviton from observation on Abell1689 galaxy cluster. They report
@@
This graviton mass is a factor of about $7$ lighter than our $9.5\times 10^{-29}$ eV for the Milky Way. This means that the graviton from Abell 1689 reaching earth-bound observers, has trespassed a mass of $497^{2}=$ 49 Milky Ways. Abell 1689 galaxy cluster is reported to have about 1000 galaxies, [40]. A recent paper [41], employs a Yukawa cosmology to constrain the graviton mass from observations of the Milky Way and M31. Although their approach and reasoning differs from ours, the fact both they and we speak of Yukawa in galactic and cosmic contexts, conclusions are very much the same. For instance, both studies address deduction of graviton mass from observations of rotation curves, or both maintain rotation curves should be considered as quantum manifestations of galactic matters. Their graviton mass $m_{g}\approx 1.54\times 10^{-62}$ kg $=8.64\times 10^{-27}$ eV is, however, is $91$ times larger than our, $9.5\times 10^{-29}$ eV for the Milky Way.
5 Conclusion
We argue that matter in classical physics, including GR, is a collection of discrete point-like entities, to which in defiance of Heisenberg’s principle of uncertainty, one assigns precise coordinates and momenta. Singularities and divergences encountered in Schwarzschild’s solution or elsewhere in GR stem from this non-quantum act.
We propose to replace the discrete collection of point-like masses of GR by a collection of wave-like entities, KG and/or YKG waves. Waves, on the one hand, are spatially distributed entities and do not end up in spacetime singularities, as one approaches the origin. Waves, on the other hand, reach infinity and produce $\frac{1}{r}$ forces. The latter feature alone is capable of explaining the flat rotation curves of spiral galaxies, two birds with one stone.
The far zone solutions are the highlight of this paper. YKG waves reach infinity as $\frac{1}{r}\exp(-\kappa/r)\cos(kr+\alpha)$ if degenerate, and as $\frac{1}{r}\exp(-\kappa/r)$ if non-degenerate, see (26), (28) and (29) . At intermediate distances, the gravitational force is dominantly $\frac{1}{r^{2}}$, but transits to $\frac{1}{r}$ as $r\rightarrow\infty$. The latter is responsible for the flat rotation curves of spirals and can be interpreted as an alternative gravity, a dark matter scenario, or whatever.
@@ 5 Conclusion
Nowadays quantum technologists are sending and receiving quantum messages across distances of tens and hundreds of kilometers through the science and technology of quantum entanglement. Do astronomers have sufficient evidence to claim they are observing quantum effects across galaxies through rotation curves? We think they do; we maintain that the Tully-Fisher relation is the cumulative quantum relic of some $10^{70}$ nucleons in galactic matters.
In the standard GR gravitation is mediated by massless gravitons and are known to come in two polarizations. The YKG gravitons are massive and (we will show in a forthcoming communication that) they have at least three polarizations. They can be put in one-to-one correspondence with the three Yukawa pions that mediate the residual strong interactions. There are further similarities between YKG and the residual strong interactions. For instance, both are attractive forces. Or as shown in Table (4), the mass of YKG graviton, for the gravitational interaction of nucleons, and the $d$ and $u$ quarks are nearly identical to the masses of $\pi$ mesons. Given these commonalities, one may wonder whether the YKG gravitation and, at least the residual strong interactions, are the two manifestations of a single underlying concept, the two sides of the same coin.
Acknowledgments
We thank M. Vahidinia and V. Taghiloo for their participation in formulating the physics of the model elaborated in Sec.(2). We also thank H. Firouzjahi for his insightful comments on the astrophysical aspects of the paper and F. Arash for discussions on elemental aspects of nucleons mass and their internal structure.